f(x)=-x^2+12x+4
The vertex is?
The equation of the line symmetry is x=?
The maximum/minimum of f(x) is?

Answer by user_dude2008(1862)

x=-b/2a=-12/(2*-1)=-12/(-2)=6

f(6)=-(6)^2+12(6)+4=-36+72+4=40

Vertex: (6,40)

line of symmetry: x=6

Max: 40